OpenStudy (anonymous):

solve the given system of equations graphically: y=-x and y=x+2

OpenStudy (anonymous):

solve x and y?

OpenStudy (anonymous):

i honestly have no idea this came directly from a review sheet.

OpenStudy (anonymous):

To solve graphically, means graph each line and see where they intersect.

OpenStudy (anonymous):

Ok, you have grid paper to draw?

OpenStudy (anonymous):

you can graph by draw on grid paper, and solve x,y to check it correct

OpenStudy (anonymous):

Do you want me to solve for x and y?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

y=-x, you rewrite: y+x=0 y=x+2 , you rewrite: y-x=2 y+x=0 Add : y-x=2 ---------------- 2x=0 x =0/2 x = 0 subtitution x=0 find y y =0+2 y =2 (0,2) you can check by use the grid paper graph 2 line if they intersect at (0,2) then you correct ,if not maybe the graph you draw it wrong or solve eqaution it wrong,I hope it help

OpenStudy (anonymous):

sorry I tye wrong my keybroad not work propertly, equation not eqaution

OpenStudy (anonymous):

Am I nuts? y+x=0 y-x=2 --------------add 2y+0=2 therefore y=1 x=-1

OpenStudy (anonymous):

Graphically, y=-x is a -45 degree line that crosses the axes at the origin (0,0) and y=x+2 is a +45 degree line that crosses the Y axis at 2 (if x=0, y=2) the two lines will cross at (-1,1)

OpenStudy (anonymous):

sorry 2y but I writ 2x alanhuth is right